Opened 9 years ago

Closed 7 years ago

Last modified 3 years ago

#358 closed flight processing (fixed)

EU10/03 & EUFAR10/01 (AIMWETLAB), flight day 229b/2010, Kis Balaton

Reported by: iopa Owned by: knpa
Priority: alpha 5 Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors: emca

Description (last modified by emca)

Data location: ~arsf/workspace/EU10_03-2010_229b_Kis_Balaton

Data arrived from ARSF via SATA disk Q on 15/09/2010.

Scientific objective: Lake Balaton provides an ideal area for testing emerging algorithms on retrieval of in-water constituents and on mapping bloom-forming cyanobacteria using in-situ radiometric, airborne and satellite data.

Priority: a5 - the data set as a whole is listed as a5 due to it being part of EU10/03, however the PI of EU10/03 has lowered the LiDAR priority to a4m, and therefore LiDAR stands with a priority of a4h due to being part of EUFAR10/01.

PI: A. Tyler (EU10/03) & A. Zlinszky (EUFAR10/01)

When data is processed it should be dispatched to both PIs individually for their respective projects.

Sensors:

  • Eagle (20/10/2010 & 28/10/2010)
  • Hawk (28/01/2011)
  • Leica LIDAR (28/03/2011 & 19/04/2011)
  • Digital Photography (20/10/2010 & 28/10/2010)

Change History (45)

comment:1 Changed 9 years ago by iopa

  • Description modified (diff)

comment:2 Changed 9 years ago by anch

  • Description modified (diff)

Note: Data is currently symlinked from ~arsf/arsf_data/2010/flight_data/archive_overflow and when this is resolved the symlinks will need to be re-directed.

comment:3 Changed 9 years ago by anch

Applanix and IPAS processing for this project are complete

comment:4 Changed 9 years ago by anch

  • Owner set to anch
  • Status changed from new to accepted

Processing hyperspectral for this project now.

comment:5 Changed 9 years ago by iopa

  • Description modified (diff)

comment:6 Changed 9 years ago by anch

Camera processing for this project is ready for delivery check.

comment:7 Changed 9 years ago by anch

Eagle is ready for delivery check

comment:8 Changed 9 years ago by emca

Delivery checked Photography data
~airborne/workspace/EU10_03-2010_229b_Kis_Balaton/delivery/20101007/EU10_03
No problems - ready to deliver

Delivery checked Hyperspectral data
~airborne/workspace/EU10_03-2010_229b_Kis_Balaton/delivery/20101006/EU10_03

All processed files need to have flight letter added (229b - lev1 files & jpgs)
Missing first flight line - either process & include this line or explain its absence in the readme. If it is not going to be delivered note this in the ticket.
Add DEM to delivery
Change permissions of *.bil files so they are readable to all (as this will cause problems when copying to disk)
Misc directory is empty - add appropriate separation file here

Readme:
Add flight letter (229b) to each filename
Add DEM to delivery contents
DEM Section - "this text is subject to change" - please remove & update ASTER DEM information
Azgcorr example command - change DEM name to match delivered ASTER DEM

comment:9 Changed 9 years ago by emca

Also need to add Eagle SCT offsets to the ticket

comment:10 Changed 9 years ago by emca

Eagle and Camera data send to A Tyler 20/10/2010.

Will also need to be sent to A. Zlinszky

comment:11 Changed 9 years ago by anch

SCT Offsets for Eagle lines are:

EAGLE
-1 : 0.03
-2 : 0.05
-3 : 0.00
-4 : 0.00
-5 : 0.02
-6 : 0.02
-7 : 0.01
-8 : 0.02
-9 : 0.03
-10 : 0.03
-11 : 0.00
-12 : 0.00
-13 : 0.02
-14 : 0.05
-15 : 0.03
-16 : 0.93
-17 : 0.03

Flightline -1 was aborted due to cloud as mentioned in the log sheet, however I still found an SCT for it (Having not realised it was aborted) and I've listed that here.

comment:12 Changed 9 years ago by anch

Decided to deliver flightline 1 afterall as there was a fair amount of data and it may be useful

comment:13 Changed 9 years ago by knpa

First 12 bands of Hawk lev1s are blank. The raws are fine. Appears to be an issue with the bandset labels in the header files, they start at 900nm which is lower than it should be. Possibly incorrect bandset selected or none specified. Will leave this until end of the year and new Hawk calibration.

comment:14 Changed 9 years ago by emca

  • Description modified (diff)

comment:15 Changed 9 years ago by iopa

  • Description modified (diff)

Dispatched Eagle and Camera to Dr. Zlinszky on 28/10/2010.

comment:16 Changed 8 years ago by anch

Starting Hawk processing for this project

comment:17 Changed 8 years ago by anch

Hawk ready for delivery checking

comment:18 Changed 8 years ago by anch

SCTs for Hawk - lots of zeros!

HAWK
-1 : 0.00
-2 : 0.00
-3 : -0.02
-4 : 0.00
-5 : 0.00
-6 : 0.00
-7 : 0.00
-8 : 0.00
-9 : 0.00
-10 : 0.00
-11 : 0.00
-12 : 0.00
-13 : 0.00
-14 : 0.00
-15 : 0.00
-16 : 0.29
-17 : 0.00

comment:19 Changed 8 years ago by knpa

Starting Hawk delivery check.

comment:20 Changed 8 years ago by knpa

The hawk data had to be reprocessed (twice?) because of the mistakes with the hawk calibration.

Restarting hawk delivery check.

comment:21 Changed 8 years ago by knpa

Delivery check complete.

All looks fine apart from:

  1. Typo in Readme: "Some cloud on line 22 (Filename h229b051b.bil) has causeD overflows"
  2. Need to mention regular dropped scan lines in data quality section
  3. Need to change permissions of level 1 files to read all.

Remember that this needs to be dispatched to two different PIs, with alterations to Readme file.

comment:22 Changed 8 years ago by anch

Hawk delivered 28/01/2011

comment:23 Changed 8 years ago by emca

  • Other processors set to emca

Starting LiDAR processing

comment:24 Changed 8 years ago by anch

LiDAR sent to A. Zlinszky at his request before roll correction/classification

comment:25 Changed 8 years ago by emca

LiDAR is ready for delivery check at
~airborne/workspace/EU10_03-2010_229b_Kis_Balaton/delivery/20110310

comment:26 Changed 8 years ago by knpa

Taking LIDAR delivery check.

comment:27 Changed 8 years ago by knpa

Delivery check complete.

No problems found - ready to deliver.

Remember that this project needs to be delivered to Dr Zlinszky as well as Dr Tyler.

comment:28 Changed 8 years ago by anch

  • Summary changed from EU10/03 & EUFAR10/01, flight day 229b/2010, Kis Balaton to EU10/03 & EUFAR10/01 (AIMWETLAB), flight day 229b/2010, Kis Balaton

comment:29 Changed 8 years ago by iopa

  • Description modified (diff)

LiDAR data dispatched to Dr Zlinszky. (28/03/2011)

comment:30 Changed 8 years ago by emca

  • Description modified (diff)

LiDAR data sent to A Tyler 19/04/2011

comment:31 Changed 8 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from anch to knpa
  • Status changed from accepted to assigned

Project has been r-synced. Beginning archiving.

comment:32 Changed 8 years ago by knpa

Tabi data present.

AIMMS and GRIMM data present.

comment:33 Changed 8 years ago by knpa

Project has been tarballed and made available to NEODC (12/05/11).

comment:34 Changed 7 years ago by knpa

Backed-up by NEODC and moved to archive disk.

comment:35 Changed 7 years ago by emca

Reprocessed data using APL software

No sync for Eagle line 16, sct value of 0.935 used to process.

comment:36 Changed 7 years ago by knpa

  • Milestone set to 2010 data processing completion

Removed non-reprocessing workspace version. Leaving ticket open for reprocessing.

comment:37 Changed 7 years ago by emca

Ho sync for hawk line 16 either - sct value of 0.235 used.

Hyperspectral data now ready for delivery check

comment:38 Changed 7 years ago by knpa

Delivery checking the reprocessed hyperspectral.

comment:39 Changed 7 years ago by knpa

Mapped files look OK - zipping them up as we urgently need the space.

comment:40 Changed 7 years ago by knpa

Delivery check complete.

All looks good apart from:

  • Need to mention the regular dropped scan lines in Hawk
  • Logsheet name incorrect: should have - instead of _ before site name.

comment:41 Changed 7 years ago by emca

Updated delivery for above comments.

Reprocessing complete. Ready for delivery.

comment:42 Changed 7 years ago by emca

Data sent to Dr Zlinszky 16/08/2012

comment:43 Changed 7 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

comment:44 Changed 3 years ago by dac

Archiving

Processed data available from NEODC (http://browse.ceda.ac.uk/browse/neodc/arsf/2010/EU10_03/EU10_03-2010_229b_Kis_Balaton) started uploading raw.

comment:45 Changed 3 years ago by dac

Archiving

Raw data also available from NEODC.

Note: See TracTickets for help on using tickets.